OpenStudy (anonymous):

Find the equation of the line that passes through (−2,5) and is perpendicular to the line 2x + y = 5 . Write your answer in slope-intercept form.

OpenStudy (anonymous):

if a line is perpedicular to another line, the product of their slopes will equals to -1 suppose m is the slope. from 2x + y = 5 -> y=-2x+5 , we get m1= -2 m1*m2 = -1 (-2) * m2 =-1 -> m2= 1/2 so the slope of the other line is 1/2 then the equation of the other line will be y=(1/2)x +b we know that the line passes through (−2,5) so -> 5=(1/2)(-2) + b 5=-1+b -> b=6 hence the equation of the line is y=(1/2)x + 6